# Data Augmentation for Traffic Classification

### 19 Jan 2024 v1

@@ 6 Discussion and Conclusions
In this work we presented a benchmark of hand-crafted DA for TC covering multiple dimensions: a total of 18 augmentations across 3 families, with 3 policies for introducing augmentations during training, investigating the classification performance sensitivity with respect to augmentations magnitude and class-weighted sampling across 3 datasets with different sizes and number of classes. Overall, our results confirm what previously observed in CV literature—augmentations are beneficial even for large datasets, but in absolute terms the gains are dataset-dependent. While from a qualitative standpoint, sequence and mask augmentations better suit a TC tasks compared to amplitude augmentations, no single augmentation is found superior to alternatives and combining them (via stacking or ensembling), even when selecting top-performing ones, does not improve performance compared to using augmentations in isolation.
By investigating the latent space geometry of models trained with augmentations, we confirm that effective augmentations provide good sample variety by creating samples neither too similar nor too different from the original samples that foster the extraction of better data representations by making model training more complex (as reflected by longer training times).
Despite the multiple dimensions covered, our work suffers from some limitations. Most notably, it would be desirable to include the larger and recent $\mathtt{CESNET\text{-}TLS}22$ [26] and $\mathtt{CESNET\text{-}QUIC}23$ [24] datasets but such expansion requires large computational power.121212 For reference, models trained on $\mathtt{Enterprise}$ can take up to 6 hours. Since CESNET datasets contains 100$\times$ the number of samples of $\mathtt{Enterprise}$, performing a thorough exploration of the DA design space is extremely resource demanding. Still related to using large datasets, we can also envision more experiments tailored to investigate the relationship between datasets size and augmentations. For instance, one could sample down (e.g., randomly select 1% or 10% of the available samples) a large dataset and investigate if augmentations result more effective in the reduced datasets. In particular, since Inject shows a positive trend with respect to its intensity $N_{inject}$ we hypothesize that by augmenting a small dataset one can achieve the same performance as using larger datasets—showing these effect is clearly relevant for TC as collecting and releasing large datasets is currently a pain point. Last, our modeling campaign consider a single model architecture but benchmarking DA via alternative architecture, e.g., Transformers-based for time series [46].
@@
Overall, we believe that the performance observed in our experimental campaign might still represent a lower bound and extra performance improvements could be achieved via generative models. We call for the research community to join us in our quest for integrating DA and improve TC performance.

### 23 Jan 2024 v2

@@ 6 Discussion and Conclusions
In this work we presented a benchmark of hand-crafted DA for TC covering multiple dimensions: a total of 18 augmentations across 3 families, with 3 policies for introducing augmentations during training, investigating the classification performance sensitivity with respect to augmentations magnitude and class-weighted sampling across 3 datasets with different sizes and number of classes. Overall, our results confirm what previously observed in CV literature—augmentations are beneficial even for large datasets, but in absolute terms the gains are dataset-dependent. While from a qualitative standpoint, sequence and mask augmentations are better suited for TC tasks than amplitude augmentations, no single augmentation is found superior to alternatives and combining them (via stacking or ensembling), even when selecting top-performing ones, marginally improves performance compared to using augmentations in isolation. Last, by investigating the models latent space geometry, we confirm that effective augmentations provide good sample variety by creating samples that are neither too similar nor too different from the original ones which fosters better data representations extraction (as suggested by the longer training time).
Despite the multiple dimensions covered, our work suffers from some limitations. Most notably, it would be desirable to include the larger and more recent $\mathtt{CESNET\text{-}TLS}22$ [25] and $\mathtt{CESNET\text{-}QUIC}23$ [24] datasets but such expansion requires large computational power.121212 For reference, models trained on $\mathtt{Enterprise}$ can take up to 6 hours. Since CESNET datasets contains 100$\times$ the number of samples of $\mathtt{Enterprise}$, performing a thorough exploration of the DA design space is extremely resource demanding. Still related to using large datasets, we can also envision more experiments tailored to investigate the relationship between datasets size and augmentations. For instance, one could sample down a large dataset (e.g., by randomly selecting 1% or 10% of the available samples) and investigate if augmentations result more effective with the reduced datasets. In particular, since Inject shows a positive trend with respect to its intensity $N_{inject}$ we hypothesize that by augmenting a small dataset one can achieve the same performance as using larger datasets—showing these effects are clearly relevant for TC as collecting and releasing large datasets is currently a pain point. Last, our campaigns rely only a CNN-based architecture while assessing DA with other architectures (e.g., Transformer-based for time series [45]) is also relevant.
@@
Overall, we believe that the performance observed in our experimental campaigns might still represent a lower bound and extra performance improvements could be achieved via generative models. We call for the research community to join us in our quest for integrating DA and improve TC performance.